OpenStudy (honeymonster):

again, this equation makes no sense to me :/ please any help would be very much appreciated..posted in replies

OpenStudy (honeymonster):

\[\frac{ x+1 }{ 3x }=\frac{ x+2 }{ 3x+1 }\]

OpenStudy (anonymous):



OpenStudy (honeymonster):

thank you sooo much :)..

OpenStudy (anonymous):

anytime